OpenStudy (lukecrayonz):

For the function, find the pts on the graph at which the tangent line is horiziontal. If none exist, state that fact. f(x)=5x^2-2x+5

OpenStudy (anonymous):

Do you know what a derivative is?

OpenStudy (lukecrayonz):

Yes

OpenStudy (anonymous):

lines are horizontal when the slope is 0

OpenStudy (lukecrayonz):

So just find the derivative then set it equal to zero?

OpenStudy (anonymous):

yes

OpenStudy (lukecrayonz):

Gotchya:) Makes sense

OpenStudy (lukecrayonz):

Wait, question :o

OpenStudy (lukecrayonz):

@wio

OpenStudy (lukecrayonz):

So the deriv is y=10x-2

OpenStudy (anonymous):

yes

OpenStudy (lukecrayonz):

0=10x-2 +2 2=10x 2/10=.2 .2=x

OpenStudy (lukecrayonz):

So my answer is just (0.2,0)?

OpenStudy (lukecrayonz):

I feel like there has to be another point :O

OpenStudy (anonymous):

Why, it's a parabola.

OpenStudy (anonymous):



OpenStudy (lukecrayonz):

Well the math program I'm using is saying it's wrong too :(

OpenStudy (anonymous):

where else is it horizontal?

OpenStudy (lukecrayonz):

It says the answer is (1/5,24/5)

OpenStudy (anonymous):

0.2 =1/5

OpenStudy (lukecrayonz):

That's just the vertex.. how in the world did they get 24/5

OpenStudy (lukecrayonz):

I'm assuming the answer to these questions will always just be the vertex?

OpenStudy (lukecrayonz):

Yeah, that's just dumb. The answers to these are the vertex.